OpenStudy (needhelpp101):

Medal and Become Fan Rewarded for Correct Information/Answer... Please help me :D Find the odds. The kings are separated from a deck of standard playing cards and shuffled. One king is randomly selected. What are the odds in favor of drawing a black card? A. 4 to 1 B. 1 to 4 C. 1 to 2 D. 1 to 1

OpenStudy (kropot72):

There are 4 kings in a standard pack. Two of the 4 kings are in black suits. The odds are the ratio of the probability of drawing a black card to the probability of not drawing a black card. Can you see what the correct choice is?

OpenStudy (kropot72):

No. What is the probability of drawing a black card?

OpenStudy (needhelpp101):

A. 4 to 1

OpenStudy (needhelpp101):

I meant B

OpenStudy (kropot72):

Can you please bear with me while I take you through the steps. There are 4 cards, 2 of which are black. So what is the probability of drawing a black card?

OpenStudy (needhelpp101):

There are 2 ways you can draw a black king from the 4 possibilities. 2/4 = 1/2 = 50%

OpenStudy (kropot72):

Good work! You are correct. There are 4 cards, 2 of which are not black. So what is the probability of drawing a card that is not black?

OpenStudy (needhelpp101):

thx but i'm still a little confused..I dont know why :(

OpenStudy (kropot72):

Can you just answer this and then the final answer will soon emerge: There are 4 cards, 2 of which are not black. So what is the probability of drawing a card that is not black?

OpenStudy (kropot72):

Do you think that the probabilities of drawing a card that is black and a card that is not black are the same?

OpenStudy (needhelpp101):

1 to 2

OpenStudy (kropot72):

The odds are the ratio of the probability of drawing a black card to the probability of not drawing a black card. The two probabilities are equal in this case. Can you now see what the correct choice of the odds must be?

OpenStudy (needhelpp101):

yeah Thx :D

OpenStudy (kropot72):

You are correct! Well done :)

OpenStudy (needhelpp101):

Awesome :D

OpenStudy (kropot72):

You're welcome :)